OpenStudy (anonymous):

The graph of the following system of equations is −2x + y = 3 4x + 2y = 2 overlapping lines parallel lines intersecting lines

OpenStudy (mrnood):

first rewrite both equations so they are each in the form y=mx + b you can easily tell the answer then (I will help if oyu write the equations in that form...)

OpenStudy (anonymous):

im not sure how to

OpenStudy (mrnood):

in the first equation add 2x to each side do that one first

OpenStudy (anonymous):

am i supposed to solve it? @MrNood

OpenStudy (mrnood):

just add 2x to each side and write the resulting equation here

OpenStudy (anonymous):

−2x + y = 3 2x 2x

OpenStudy (mrnood):

In that case I'm afraid you don't have the background skills to answer a question like this. You should go back to study simple equations, and how to re-arrange them, and then come back to these 'systems of equations'

OpenStudy (anonymous):

@imqwerty can you explain this pls

OpenStudy (mrnood):

no one can 'explain' this in time for you to do an exam. you need to re-arrange the 2 equations so that they look lik e y=mx+b get the y term on one side, and the other terms on the other side you must have done this in your class if it being set in an exam (and btw - why are you asking exam questions here?)

OpenStudy (anonymous):

um its online class .. tyvm i basically have to teach myself gtfo

OpenStudy (anonymous):

i can take my time on virtual school on a exam its at your own pace carry on

OpenStudy (mrnood):

in the first equation add 2x to each side do that one first

imqwerty (imqwerty):

ok so both the equation represent 2 lines the standard equation of a line is--->ax+by+c=0 suppose we have 2 lines say-\[a_{1}x+b_{1}y+c_{1}=0\]and\[a_{2}x+b_{2}y+c_{2}=0\] the lines are overlapping if-\[\frac{ a_{1} }{ a_{2} }=\frac{ b_{1} }{ b_{2}}=\frac{ c_{1} }{ c_{2}}\] the lines are parallel if-\[\frac{ a_{1} }{ a_{2} }=\frac{ b_{1} }{ b_{2}} \neq \frac{ c_{1} }{ c_{2}}\] and intersecting if-\[\frac{ a_{1} }{ a_{2} }\neq\frac{ b_{1} }{ b_{2}} \neq\frac{ c_{1} }{ c_{2}}\] so we have 2 equations that is we got 2 lines i'll jst give u some hint then u can do it ..if u get any prblm jst ask :) so lets take the 1st line -2x+y=3 we need to convert it to this form-->ax+by+c=0 so basically we want 0 on the right so we subtract 3 from both sides -2x+y-3=3-3 -2x+(1)y-3=0 comparing this with ax+by+c=0 we get a=-2 b=1 and c=-3 now we have a,b,c of 1st line find a,b,c of second line and then apply the thing i wrote up there^ :) which ever satisfies is ur answer

OpenStudy (mrnood):

yeah - like she'll understand that if she won't add 2x to each side.....

OpenStudy (anonymous):

so it would be intersecting right? @imqwerty

imqwerty (imqwerty):

yes :)

OpenStudy (anonymous):

Thank You :)

imqwerty (imqwerty):

no prblm :)